QUESTION 27 Q27. A manager at a local bank analyzed the relationship between monthly salary (y, in $) and length of service (QUESTION 28 Q28. An real estate analyst believes that the three main factors that influence an apartments rent in a collegeQUESTION 29 Q29. A manager at a local bank analyzed the relationship between monthly salary by, in $) and length of service (QUESTION 31 Q31. Given the following portion of regression results, which of the following conclusions is true with regard toQUESTION 33 Q33. When testing linear restrictions, the test statistic is assumed to follow the y = Bo + B1x1 +B2x2 + ε, you w
Answer #1

27)

Y = 784.92 + (9.19*48) = 1226.04

D option $1226.

28)

Correct option is b

An additional bedroom increases rent , on average by $226.

29) the standard error of the estimate is the square root of the average squared deviation.

Correct option is c 269.63

31) a option .at least one of the explanatory variable is significantly related to response variable.

32) 3781/470 = 8.04

Correct option is a. 8.04

33) as p value 0.009 is less than 0.10 so we reject null hypothesis.

Correct option is d.

Reject null hypothesis and conclude that influence of x1 and X2 on y is significant.